# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

### v3

@@ 6 Conclusion
In this paper, we revisited existing FL backdoor detection methods and identified two fundamental limitations: non-IID divergence and out-of-distribution (OOD) prediction bias. To tackle these challenges, we designed a new method, Coward, which leverages a collision-based OOD watermark that the server injects and later inspects to expose malicious clients. Inspired by the multi-backdoor collision effects, We carefully design Coward to enable an inverted detection paradigm and to minimally interfere with benign pattern learning, thereby mitigating the severity of OOD prediction bias and its adverse effects on misjudgment. Extensive experiments verified the effectiveness of Coward and its resilience against potential adaptive attacks.
(a) w./ BN swith
@@
Figure 2: OOD watermark collision under centralized scenario. Our OOD watermark is planted as the second backdoor. The resulting collision effect is significant, regardless of whether the BN layer is switched. However, switching the BN layer creates a more pronounced performance discrepancy between benign finetuning and backdoor injection.
@@
Appendix A Collision Effects on OOD Watermark
In Section 3.4, we introduced the multi-backdoor collision effect under a centralized setting where two backdoors share the same dataset, aiming to illustrate its core mechanism. To further demonstrate how this effect manifests in our OOD-based watermark under practical FL scenarios, we provide an extended analysis in this section. Specifically, we examine the collision effect between the attacker’s backdoor and our OOD watermark across progressively complex scenarios: starting from a centralized setup, moving to a static FL environment with fixed client participation, and finally to the dynamic FL scenario adopted in our main experiments. To align with main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default. As elaborated in the following sections, results across all three stages consistently reveal a clear collision effect, validating the capability of our OOD watermark to trigger such interactions and effectively distinguish malicious clients.
Centralized Scenario. We conduct a similar experiment as in Section 3.4, with the key difference that the second backdoor injection is replaced by our OOD watermark planting process. Specifically, we use a target label 8 watermark with MNIST as the planting set. As shown in Fig. 2, in each subfigure, the left side illustrates the collision effect, while the right side shows the discrepancy between benign ($i.e.$, finetuning) and malicious ($i.e.$, backdoor injection) operation. Specifically, in subfigure (a), we observe a clear collision effect on the OOD watermark, characterized by a clear drop in ASR and a notable distinction between benign and malicious clients. In subfigure (b), when the BN layer switch is disabled, the benign-malicious discrepancy becomes indistinguishable, indicating the necessity of BN isolation for effective detection.
Static FL Scenario. We consider a slightly more complex scenario under a static FL setting, involving 10 clients with full participation in each communication round. Among them, only one client is malicious. This simplified full-participation setup facilitates the observation of each client’s temporal behavior and helps illustrate the collision effect in a federated learning context. In each global round, the server injects an OOD watermark, while all benign clients perform standard local training.To clearly observe the collision effect, the attacker is configured to behave benignly for the first 20 rounds and initiates a backdoor attack targeting label 0 from round 21 onward. After each local update, we inspect the watermark accuracy of all 10 participated clients. As shown in Figure 3(b), once the malicious client begins the attack, its watermark accuracy drops sharply to nearly zero, indicating a strong collision effect. In contrast, the remaining nine benign clients consistently maintain high watermark accuracy. This sharp divergence highlights the benign-malicious discrepancy induced by the collision effect, which is critical for enabling effective backdoor detection. Additionally, in subfigure (a), we observe that benign clients struggle to preserve the watermark without BN layer isolation, underscoring the importance of BN switching in reducing FPR.
Dynamic FL Scenario. We finally present the collision effect under the dynamic FL setting used in our main experiments, where 10 out of 100 clients are randomly selected to participate in each global round. This dynamic participation leads to divergent and inconsistent watermark retention behaviors among benign clients, making direct comparisons between them infeasible. Therefore, we focus on observing relative trends across rounds. Specifically, we plot the temporal variation of watermark accuracy for all clients based on their actual participation order. To control page length, we select clients at intervals of 3 ($i.e.$, client IDs with a gap of 3). As shown in Figure 1, we observe the following: (1) Malicious client 0 exhibits a pronounced collision effect, with watermark accuracy dropping to nearly zero. (2) Benign clients maintain relatively high watermark accuracy (mostly above 50%), though the magnitude and variation patterns differ across clients due to heterogeneous participation and client distribution. Overall, the results reveal a clear separation between benign and malicious clients, enabling robust and tolerant threshold selection. Despite being influenced by dynamic participation and non-i.i.d.data, the collision effect is still effective in distinguishing attackers.
TABLE I: Performance of Coward with different OOD Base Mapping on the CIFAR-10 dataset.
@@ Appendix A Collision Effects on OOD Watermark
Default 100.0 2.4 9.8 91.3
Shift+1 99.5 3.0 9.8 91.5
Shift+3 99.5 2.9 9.8 91.5
Shift+5 100.0 2.6 9.7 91.3
TABLE II: Performance of Coward for all watermark labels.
@@ Appendix B Discussion on Watermark Configuration
Our detection paradigm is flexible with respect to specific watermark configurations, allowing practical deployment scenarios to select or design their own. To concretely demonstrate this flexibility, we provide empirical evidence by evaluating constituent components of our OOD watermark: the OOD base mapping, the watermark trigger, and the watermark target. To align with main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default.
OOD Base Mapping. The key role of the OOD base mapping is to provide a fixed foundation for building the watermark mapping. As long as the mapping remains class-wise and fixed, the specific choice of mapping is not critical. While we adopt a diagonal mapping in the main experiments for simplicity, we demonstrate here that our approach is compatible with other mapping configurations, highlighting its flexibility. Specifically We replace the default diagonal mapping ($i.e.$, $i\rightarrow i$) with three alternative 1-to-1 shift-based mappings: $i\rightarrow(i{+}3)\bmod 10$, $i\rightarrow(i{+}5)\bmod 10$, and $i\rightarrow(i{+}7)\bmod 10$. As shown in Table I, all variants achieve consistently strong detection performance, indicating that Coward does not rely on specific OOD mapping rule. These results confirm that our method is robust to the choice of OOD assignment and can be seamlessly adapted to different mapping schemes.
Trigger Type. Previously, we have already presented quantitative results demonstrating the general effectiveness of the collision effect and the corresponding detection mechanism across various trigger pairs in Table VI. To provide a more intuitive understanding, we further visualize the appearance of the triggers. As illustrated in Figure 4, our server-malicious trigger pairs include both semantically similar combinations ($e.g.$, a server-side square pattern and a client-side pixel-style trigger) and semantically dissimilar ones ($e.g.$, a server-side WaNet trigger and a client-side semantic trigger). Notably, both types consistently yield strong detection performance. These results indicate that the effectiveness of our approach is largely independent of the visual semantics of the triggers, offering defenders a broad design space to customize trigger appearances according to their specific needs.
Watermark Label. In the main experiments, we fixed the watermark target label to 8 for simplicity. However, we emphasize that the effectiveness of Coward is generally applicable to any target label. To validate this flexibility, we present results in Table II, where the attack target is fixed to label 0 and the watermark label is varied across all possible values. The consistently strong detection performance across different watermark labels confirms the adaptability of our approach in selecting the watermark target.
TABLE III: Ablation study of Coward on watermark injection strength.
@@
TABLE VII: Performance of Coward under different levels of global model regularization.
@@ Appendix C Ablation Study on Watermark Injection
Beyond the design choices of the watermark itself, an equally important yet distinct problem is how to effectively implant the watermark into the global model. In this section, we conduct ablation studies on all factors that influence the implantation process. To align with our main experiments, we conduct all evaluations under the 0.9 non-IID setting on the CIFAR-10 dataset by default.
Impact of Global Model Regularization. Table VII reports the effect of varying levels of global model regularization. We observe a positive impact on TPR, where applying regularization consistently yields higher TPRs compared to the no-regularization baseline. Increasing the regularization strength $\lambda$ leads to slight improvements in TPR, with the best performance observed at $\lambda=0.3$. These results suggest that a moderate level of regularization is beneficial for achieving optimal detection performance.
Learning Efficiency of Watermark Injection. We conduct ablation studies on three key hyperparameters that control the strength of watermark injection: the number of OOD watermark samples, the watermark learning rate, and the number of injection iterations per global round. We vary one factor at a time while keeping the others fixed. By default, the watermark learning rate is set to 0.001, the number of watermark injection rounds to 5, and the watermark dataset size to 1000. As shown in Table VI, Coward consistently achieves strong performance, with TPR exceeding 98% and FPR remaining below 10% in most cases. Even under extremely constrained injection settings ($e.g.$, a watermark dataset size of 200, a learning rate of 0.0001, or a single injection round), the method still maintains competitive results, achieving 100% TPR with FPR below 25%. Overall, effective watermark injection requires only a low level of effort. Most parameter variations lead to only marginal performance differences, and any configuration beyond the most extreme cases yields satisfactory performance. These results indicate that our method is robust across a wide range of watermark injection strengths and introduces negligible computational overhead in federated learning. Notably, effective defense can be achieved with as few as only 1 round of watermark injection.
Appendix D Potential Limitations and Future Directions
While our method provides a simple and effective early solution for proactive defense, there remain several open directions that worth further exploration and refinement in future works on this topic:
@@ Appendix D Potential Limitations and Future Directions
  • Multi-Metric and Multi-Mechanism Detection. While our method have already achieved effective detection using a single metric ($i.e.$, watermark accuracy) with clear and consistent signals, the broader flexibility of the proposed framework allows for further expansion. Specifically, it can support joint observation of client responses to multiple operations, opening up the possibility of incorporating diverse detection mechanisms. For example, one may simultaneously leverage the maintenance and collision effects by separately monitoring OOD mappings and watermark mappings. This enriched perspective could provide complementary signals and represents a promising future direction.

### v4

@@ 6 Conclusion
In this paper, we revisited existing FL backdoor detection methods and identified two fundamental limitations: non-IID divergence and OOD prediction bias. To tackle these challenges, we designed a new method, Coward, which leverages a collision-based OOD watermark that the server injects and later inspects to expose malicious clients. Inspired by the multi-backdoor collision effects, we carefully designed Coward to enable an inverted detection paradigm and to minimally interfere with benign pattern learning, thereby mitigating the severity of OOD prediction bias and its adverse effects on misjudgment. Extensive experiments verified the effectiveness of Coward and its resilience against potential adaptive attacks.
Acknowledgement
We sincerely thank Dr. Kangjie Chen and Prof. Dacheng Tao from Nanyang Technological University for their constructive suggestions on early drafts of this work.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used ChatGPT in order to improve language clarity and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the article.
Supplementary
Appendix A Collision Effects on OOD Watermark
In Section 3.4, we introduced the multi-backdoor collision effect under a centralized setting where two backdoors share the same dataset, aiming to illustrate its core mechanism. To further demonstrate how this effect manifests in our OOD-based watermark under practical FL scenarios, we provide an extended analysis in this section. Specifically, we examine the collision effect between the attacker’s backdoor and our OOD watermark across progressively complex scenarios: starting from a centralized setup, moving to a static FL environment with fixed client participation, and finally to the dynamic FL scenario adopted in our main experiments. To align with the main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default. As elaborated in the following sections, results across all three stages consistently reveal a clear collision effect, validating the capability of our OOD watermark to trigger such interactions and effectively distinguish malicious clients.
(a) w./ BN switch
@@ Appendix A Collision Effects on OOD Watermark
Figure 1: OOD watermark collision under centralized scenario. Our OOD watermark is planted as the second backdoor. The resulting collision effect is significant, regardless of whether the BN layer is switched. However, switching the BN layer creates a more pronounced performance discrepancy between benign fine-tuning and backdoor injection.
Centralized Scenario. We conduct a similar experiment as in Section 3.4, with the key difference that the second backdoor injection is replaced by our OOD watermark planting process. Specifically, we use a target label 8 watermark with MNIST as the planting set. As shown in Fig. 1, in each subfigure, the left side illustrates the collision effect, while the right side shows the discrepancy between benign ($i.e.$, fine-tuning) and malicious ($i.e.$, backdoor injection) operations. Specifically, in subfigure (a), we observe a clear collision effect on the OOD watermark, characterized by a clear drop in ASR and a notable distinction between benign and malicious clients. In subfigure (b), when the BN layer switch is disabled, the benign-malicious discrepancy becomes indistinguishable, indicating the necessity of BN isolation for effective detection.
@@
Static FL Scenario. We consider a slightly more complex scenario under a static FL setting, involving 10 clients with full participation in each communication round. Among them, only one client is malicious. This simplified full-participation setup facilitates the observation of each client’s temporal behavior and helps illustrate the collision effect in a federated learning context. In each global round, the server injects an OOD watermark, while all benign clients perform standard local training. To clearly observe the collision effect, the attacker is configured to behave benignly for the first 20 rounds and initiates a backdoor attack targeting label 0 from round 21 onward. After each local update, we inspect the watermark accuracy of all 10 participating clients. As shown in Figure 2(b), once the malicious client begins the attack, its watermark accuracy drops sharply to nearly zero, indicating a strong collision effect. In contrast, the remaining nine benign clients consistently maintain high watermark accuracy. This sharp divergence highlights the benign-malicious discrepancy induced by the collision effect, which is critical for enabling effective backdoor detection. Additionally, in subfigure (a), we observe that benign clients struggle to preserve the watermark without BN layer isolation, underscoring the importance of BN switching in reducing FPR.
Dynamic FL Scenario. We finally present the collision effect under the dynamic FL setting used in our main experiments, where 10 out of 100 clients are randomly selected to participate in each global round. This dynamic participation leads to divergent and inconsistent watermark retention behaviors among benign clients, making direct comparisons between them infeasible. Therefore, we focus on observing relative trends across rounds. Specifically, we plot the temporal variation of watermark accuracy for all clients based on their actual participation order. To control page length, we select clients at intervals of 4 ($i.e.$, client IDs with a gap of 4). As shown in Figure 3, we observe the following: (1) Malicious client 0 exhibits a pronounced collision effect, with watermark accuracy dropping to nearly zero. (2) Benign clients maintain relatively high watermark accuracy (mostly above 50%), though the magnitude and variation patterns differ across clients due to heterogeneous participation and client distribution. Overall, the results reveal a clear separation between benign and malicious clients, enabling robust and tolerant threshold selection. Despite being influenced by dynamic participation and non-i.i.d. data, the collision effect remains effective in detection.
@@
Table 1: Performance of Coward with different OOD mappings on the CIFAR-10 dataset.
@@
Table 2: Performance of Coward for all watermark labels.
@@ Appendix B Discussion on Watermark Configuration
Our detection paradigm is flexible with respect to specific watermark configurations, allowing practical deployment scenarios to select or design their own. To concretely demonstrate this flexibility, we provide empirical evidence by evaluating constituent components of our OOD watermark: the OOD base mapping, the watermark trigger, and the watermark target. To align with the main experiments, we conduct all evaluations under the 0.9 non-i.i.d. setting on the CIFAR-10 dataset by default.
OOD Base Mapping. The key role of the OOD base mapping is to provide a fixed foundation for building the watermark mapping. As long as the mapping remains class-wise and fixed, the specific choice of mapping is not critical. While we adopt a diagonal mapping in the main experiments for simplicity, we demonstrate here that our approach is compatible with other mapping configurations, highlighting its flexibility. Specifically, we replace the default diagonal mapping ($i.e.$, $i\rightarrow i$) with three alternative 1-to-1 shift-based mappings: $i\rightarrow(i{+}3)\bmod 10$, $i\rightarrow(i{+}5)\bmod 10$, and $i\rightarrow(i{+}7)\bmod 10$. As shown in Table 1, all variants achieve consistently strong detection performance, indicating that Coward does not rely on a specific OOD mapping rule. These results confirm that our method is robust to the choice of OOD assignment and can be seamlessly adapted to different mapping schemes.
Trigger Type. Previously, we have already presented quantitative results demonstrating the general effectiveness of the collision effect and the corresponding detection mechanism across various trigger pairs in Table 5. To provide a more intuitive understanding, we further visualize the appearance of the triggers. As illustrated in Figure 4, our server-malicious trigger pairs include both semantically similar combinations ($e.g.$, a server-side square pattern and a client-side pixel-style trigger) and semantically dissimilar ones ($e.g.$, a server-side WaNet trigger and a client-side semantic trigger). Notably, both types consistently yield strong detection performance. These results indicate that the effectiveness of our approach is largely independent of the visual semantics of the triggers, offering defenders a broad design space to customize trigger appearances according to their specific needs.
Watermark Label. In the main experiments, we fixed the watermark target label to 8 for simplicity. However, we emphasize that the effectiveness of Coward is generally applicable to any target label. To validate this flexibility, we present results in Table 2, where the attack target is fixed to label 0 and the watermark label is varied across all possible values. The consistently strong detection performance across all watermark labels confirms the adaptability of our approach in selecting the watermark target.
Table 3: Ablation study of Coward on watermark injection strength.
@@
Table 4: Performance of Coward under different levels of global model regularization.
@@ Appendix C Ablation Study on Watermark Injection
Beyond the design choices of the watermark itself, an equally important yet distinct problem is how to effectively implant the watermark. In this section, we conduct ablation studies on all factors that influence watermark injection. To align with the main experiments, we conduct all evaluations under the 0.9 non-IID setting on the CIFAR-10 dataset by default.
Impact of Global Model Regularization. Table 4 reports the effect of varying levels of global model regularization. We observe a positive impact on TPR, where applying regularization consistently yields higher TPRs compared to the no-regularization baseline. Increasing the regularization strength $\lambda$ leads to slight improvements in TPR, with the best performance observed at $\lambda=0.3$. These results suggest that a moderate level of regularization is beneficial for achieving optimal detection performance.
Learning Efficiency of Watermark Injection. We conduct ablation studies on three key hyperparameters that control the strength of watermark injection: the number of OOD watermark samples, the watermark learning rate, and the number of injection iterations per global round. We vary one factor at a time while keeping the others fixed. By default, the watermark learning rate is set to 0.001, the number of watermark injection rounds is set to 5, and the watermark dataset size is set to 1000. As shown in Table 3, Coward consistently achieves strong performance, with TPR exceeding 98% and FPR remaining below 10% in most cases. Even under extremely constrained injection settings ($e.g.$, a watermark dataset size of 200, a learning rate of 0.0001, or a single injection round), the method still maintains competitive results, achieving 100% TPR with FPR below 25%. Overall, effective watermark injection requires only a low level of effort. Most parameter variations lead to only marginal performance differences, and any configuration beyond the most extreme cases yields satisfactory performance. These results indicate that our method is robust across a wide range of watermark injection strengths and introduces negligible computational overhead in federated learning. Notably, effective defense can be achieved with as few as 1 round of watermark injection.
Table 5: Attack success rate and benign accuracy against stealthy single-client backdoor attacks with advanced training strategies. Result highlighting follows the same rules as Table 1.
[…]
Table 6: Attack success rate and benign accuracy under multi-client backdoor attacks on the CIFAR-10 dataset. Result highlighting follows the same rules as Table 1.
[…]
Appendix D Additional ASR and BA Results
We provide the attack success rate (ASR) and benign accuracy (BA) results for the advanced single-attacker and multi-attacker experiments, respectively.
Appendix E Potential Limitations and Future Directions
While our method provides a simple and effective early solution for proactive defense, several open directions remain worth further exploration and refinement in future work on this topic:
@@ Appendix E Potential Limitations and Future Directions
  • Multi-Metric and Multi-Mechanism Detection. While our method has already achieved effective detection using a single metric ($i.e.$, watermark accuracy) with clear and consistent signals, the broader flexibility of the proposed framework allows for further expansion. Specifically, it can support joint observation of client responses to multiple operations, opening up the possibility of incorporating diverse detection mechanisms. For example, one may simultaneously leverage the maintenance and collision effects by separately monitoring OOD mappings and watermark mappings. This enriched perspective could provide complementary signals and represents a promising future direction.